### v1

# A 106K Multi-Topic Multilingual Conversational User Dataset with Emoticons

Abstract.
Instant messaging has become a predominant form of communication, with texts and emoticons enabling users to express emotions and ideas efficiently. Emoticons, in particular, have gained significant traction as a medium for conveying sentiments and information, leading to the growing importance of emoticon retrieval and recommendation systems. However, one of the key challenges in this area has been the absence of datasets that capture both the temporal dynamics and user-specific interactions with emoticons, limiting the progress of personalized user modeling and recommendation approaches. To address this, we introduce the emoticon dataset, a comprehensive resource that includes time-based data along with anonymous user identifiers across different conversations. As the largest publicly accessible emoticon dataset to date, it comprises 22K unique users, 370K emoticons, and 8.3M messages. The data was collected from a widely-used messaging platform across 67 conversations and 720 hours of crawling. Strict privacy and safety checks were applied to ensure the integrity of both text and image data. Spanning across 10 distinct domains, the emoticon dataset provides rich insights into temporal, multilingual, and cross-domain behaviors, which were previously unavailable in other emoticon-based datasets. Our in-depth experiments—both quantitative and qualitative—demonstrate the dataset’s potential in modeling user behavior and personalized recommendation systems, opening up new possibilities for research in personalized retrieval and conversational AI. The dataset is freely accessible  [^1] https://huggingface.co/datasets/metchee/u-sticker.
Resource, Sticker Retrieval, User Modeling
copyright: noneconference: Preprint; 2025; $c$s: Information systems Information retrievalccs: Information systems Dataset Personalization
1. Introduction
Figure 1. The colorful and vivid emoticons makes for an expressive and comfortable conversation.
Instant messaging (IM) has become an essential mode of communication, allowing users to interact efficiently. Beyond text, colorful and animated emoticons have emerged as a powerful medium of expression, enabling users to convey semantics and emotions comfortably and accurately (Ge, 2020; Constantin et al., 2019; Kariko and Anasih, 2019; Gygli and Soleymani, 2016; Jiang et al., 2023), as seen in Figure 1. Unlike general image datasets, which emphasize object recognition, emoticons carry rich semantic meaning, expressiveness, and emotional nuance, making their retrieval and recommendation fundamentally different from traditional image retrieval and recommendation tasks.
However, despite emoticons’ widespread use, personalized emoticon retrieval and recommendation remains an understudied area, primarily due to the lack of large-scale user emoticon interaction dataset. As shown in Table 1, while several emoticon dataset exist, they often lack user information, making them unable to identify the same user across different conversations. Additionally, they contain few emoticons per conversation, providing limited insight into user preferences or emoticon usage patterns. Furthermore, some datasets are not open-sourced, limiting reproducibility and progress of personalized emoticon retrieval and recommendation.
Table 1. Comparison with Emoticon Datasets. We bold the largest count among publicly available datasets and underline the second largest. Our dataset is cross-domain, large in scale and contain time and user ID for personalization.
Dataset Pub. Avail. Cross Domain Time Info. # Sticks./user # Emoticons Description
StickerTag (Wang et al., 2024) 13,571 Emoticon-tag pairs
StickerInt (Liang et al., 2024) Unknown 1,025 Dialogues with emoticons
StickerCLIP (Zhao et al., 2023) 820,000 Chinese emoticon-tag pairs
CSMSA (Ge et al., 2022) 16,000 Emoticons-tag pairs
SRS, PESRS (Gao et al., 2021, 2020) 6.82 320,168 Dialogues with emoticons
PerSRV (Chee et al., 2024) Unknown 543,098 Emoticon-query pairs
MCDSCS (Shi and Kong, 2024) 2.03 14,400 Dialogues with emoticons
SER30K (Liu et al., 2022) 30,739 Emoticon-sentiments pairs
MOD (Fei et al., 2021) 1.36 307 Dialogues with emoticons
Ours 16.90 370,222 Dialogues with emoticons
To address these limitations, we introduce a new emoticon dataset, the first large-scale dataset that includes both user information and emoticon-based conversations. Our contributions can be summarized as follows:
  • We present a new emoticon dataset, the largest emoticon dataset to date, containing 22k users, 370.2k emoticons and 8.3M conversation messages of texts and emoticons.
  • Our dataset is a multi-domain dataset that includes rich and diverse information. Covering 10 domains, it captures temporal, multilingual, and cross-domain behaviors that are not present in previous datasets.
  • Extensive quantitative and qualitative experiments demonstrate the dataset’s practical applications on user behavior analysis and modeling, personalized emoticon recommendation. It also holds potential for further research in areas such as personalized retrieval and conversational studies.
2. Related Work
We outline the current emoticon datasets and their applications. The terms sticker, emoticon, memes, gif mean approximately the same, dynamic or static multimedia that can be shared to express emotions. Herein, we use the word emoticon.
2.1. Emoticon Datasets
Due to the challenges in preparing emoticon conversation-based datasets, their availability remains limited, particularly for publicly accessible datasets. However, recent research on emoticons has gained more attention. Currently, there are only nine related datasets, with only three being publicly available. A summary of these datasets is provided in Table 1.
2.1.1. Availability
Many of these datasets are closed-source. Specifically, six datasets are not publicly accessible (Wang et al., 2024; Liang et al., 2024; Zhao et al., 2023; Ge et al., 2022; Gao et al., 2021; Chee et al., 2024), meaning they are either closed-source or require additional access that is often unattainable. This lack of open access presents a significant challenge for emoticon-based research.
2.1.2. Multi Topic Information
Furthermore, none of the aforementioned datasets contain cross-domain information, as the conversations or dialogues are not identified with specific topics. This limitation hinders the exploration of cross-domain emoticon usage, which could provide valuable insights.
2.1.3. User Information
Additionally, publicly available datasets are scarce in terms of user data, with 1.36 (Fei et al., 2021) and 2.03 (Shi and Kong, 2024) historical emoticons per user. This lack of data makes it difficult to perform user analysis and personalization-related tasks. Our dataset bridges this gap, providing the highest emoticon history per user, with an impressive 16.90 emoticons.
2.1.4. Multilingualism
Most existing datasets focus on one or two languages, limiting the ability of smaller communities to benefit from advancements in emoticon personalization. In contrast, our dataset is multilingual, allowing a broader range of communities to engage with these advancements.
As shown in Table 1 and our analysis, our dataset is the largest publicly available emoticon-user dataset. It features 370,222 emoticons, cross-domain behavior, rich contextual information, and comprehensive user and temporal data.
2.2. Emoticon Retrieval
Most previous research emphasizes the importance of data for emoticon retrieval. SRS (Gao et al., 2020), PESRS (Gao et al., 2021) require corresponding utterances, while Lao et al. (Lao et al., 2022) rely on manually labeled emotions, sentiments, and reply keywords. CKES (Chen et al., 2024) annotates each emoticon with a corresponding emotion. During emoticon creation, Hike Messager (Laddha et al., 2020) tags conversational phrases to emoticons. The reliance on data presents a significant limitation, as emoticons without associated information are excluded from consideration.
Gao et al. (Gao et al., 2020) use a convolutional emoticon encoder and self-attention dialog encoder for emoticon-utterance representations, followed by a deep interaction network and fusion network to capture dependencies and output the final matching score. The method selects the ground truth emoticon from a pool of emoticon candidates and its successor. Zhang et al. (Zhang et al., 2022) perform this on recommendation tasks. CKES (Chen et al., 2024) introduces a causal graph to explicitly identify and mitigate spurious correlations during training. The PBR (Xia et al., [n. d.]) paradigm enhances emotion comprehension through knowledge distillation, contrastive learning, and improved hard negative sampling to generate diverse and discriminative emoticon representations for better response matching in dialogues. PEGS (Zhang et al., 2024), StickerInt (Liang et al., 2024) generate emoticon information using multimodal models and selects emoticon responses, but does not consider personalization. StickerCLIP (Zhao et al., 2023) fine-tunes pretrained image encoders but does not consider personalization. In addition, many methods designed to rank emoticons from top-k candidates face a significant drawback in real-world emoticon retrieval scenarios, as they quickly become impractical when applied to larger datasets. Possibly due to the lack of dataset, most methods do not consider personalization.
2.3. Personalized Emoticon Recommendation
The absence of personalized emoticon recommendation can likely be attributed to the scarcity of emoticon-user datasets. PESRS (Gao et al., 2021) improves upon earlier work (Gao et al., 2020) by incorporating user preferences, but the dataset it relies on is not publicly accessible. This limited availability indirectly hampers the development of personalized emoticon recommendations and may impede further advancements in the field.
3. Our Dataset
Telegram (tel, [n. d.]) is a widely used open-source messaging platform, offering publicly accessible conversations. The interactions within these groups are natural, making them a valuable source for understanding human behavior, particularly in the context of personalization. As such, they present a promising dataset for emoticon-related research (Dargahi Nobari et al., 2017; Naseri and Zamani, 2019; Baumgartner et al., 2020; Blas et al., 2024; Khaund et al., 2020; La Morgia et al., 2025).
However, the raw data requires preprocessing before it can be utilized effectively. In the following sections, we outline the process of constructing the dataset. We begin by defining the criteria used in the construction process.
3.1. Construction Criteria
Given the significant implications of this dataset for future research, we adhere to the following criteria to ensure the quality and reliability of the data source
We firstly define the envision for the final user-emoticon dataset. Since emoticons are the focal point of this dataset, we would like a high presence of emoticons. Additionally, since we would like to analyze interactions across different domains, the final dataset should involve diverse topics. We would also like the dataset to be of high-quality; avoidance of spam, advertisement and groups with single person. Moreover, we noticed that previous studies overlooked linguistic variety, hence, we wish our final dataset to have multilingual coverage. Finally, to ensure a robust and generalizable dataset, we strive for sufficient dataset size.
We translate the above goals into the following dataset construction criteria,
  1. (1)
    Emoticon prevalence: Conversations must contain a significant number of emoticons.
  2. (2)
    Topic diversity: A wide range of discussion topics should be represented.
  3. (3)
    Authentic interactions: We focus on real user-to-user conversations while avoiding:
    1. (a)
      Announcement-based channels with a single speaker.
    2. (b)
      Conversations dominated by spam, advertisements, or inappropriate content.
  4. (4)
    Linguistic diversity: Conversations should represent multiple languages.
  5. (5)
    Scalability: We aim to capture as many suitable conversations as possible.
Following these criteria, we manually screened hundreds of conversation groups and ultimately selected 71 conversation groups to crawl their content.
3.2. Dataset Pre-processing
The dataset, crawled using Telethon (Exo, 2025), contains key components from user messages, including the unique conversation group ID, group name, message datetime, user ID, media types, message ID, message text, and emoticon details (ID, set ID, access hash, and mime type). Users may participate in multiple conversation groups.
3.2.1. Text Processing
As our dataset has been collected with multilingual diversity in mind, we identify the languages utilized for ease of downstream processing. We utilize the xlm-roberta-base-language-detection model (Conneau, 2019) for language identification. The model provides both the detected language and a confidence score, and we utilize a threshold of 0.99 used to ensure high reliability in language detection.
After language detection, we manually review the classification results, removing entries with fewer than 20 occurrences and those from administrative bots. The final language distribution is illustrated in Figure  2. In total, we detect 18 languages, with the top ten languages in order being English, Russian, French, Spanish, Chinese, Polish, Italian and German, Turkish, Portugese. As can be seen, emoticonU features 18 languages which is potentially viable for multilingualism emoticon-related task.
Figure 2. Distribution of detected languages in the dataset. The dataset features over 18 languages. The top ten languages are English, Russian, French, Spanish, Chinese, Polish, Italian and German, Turkish, Portugese.
3.2.2. Emoticon Processing
The downloaded emoticons come in three formats; .webp, .webm, and .tgs, we present the extension distribution in Figure 3. We convert all tgs files to the .gif extension using tgsconverter (Benau, 2025), .webp is converted to .png and no conversion is done to .webm. We break up dymamic images into frames and save a frame every second to prepare for later use.
Figure 3. Preliminary emoticons extension distribution. 65.6% of the emoticons are static and 34.4% are dyanimc.
3.3. Unsafe Text Detection and Replacement
We aim to moderate the safety of textual content. We perform this in two phases; (1) unsafe text detection and (2) offensive text replacement.
3.3.1. Unsafe Text Detection
We identify two types of unsafe text; offensive language and hyperlinks.
Offensive Language
Offensive text are content that could contain racism, vulgarities or other potentially offensive content. Since different languages require specialized tools for detecting offensive content, we focus on the ten most used languages in our dataset: English, Russian, French, Spanish, Chinese, Polish, Italian and German, Turkish, Portugese. For each, we employ well-established moderation models that have been published in peer-reviewed conferences, we describe the models in the following and the threshold values.
  • roberta-base-cold (Deng et al., 2022): Detects offensive language in Chinese. The model returns a binary label, positive values are tagged as offensive language.
  • twitter-roberta-base-offensive (Barbieri et al., 2020): Identifies offensive language in English. The model returns a 0 to 1 score for offensive and not-offensive tags. We set the offensive threshold as 0.87.
  • russian-sensitive-topic (Babakov et al., 2021): Flags sensitive or offensive content in Russian. The model returns ”sensitive” or ”not-sensitive” tags. We set ”sensitive” content as offensive.
  • toxic-bert (Davidson et al., 2019): A multilingual model that detects offensive language in French, Turkish, Portuguese, Italian, Spanish, English, and Russian. The model returns a score from 0 to 1. We set offensive threshold to be less than 0.0004.
  • dehatebert-mono-polish (Aluru et al., 2020): Detects hate speech in Polish language. The model returns a binary label, positive values are tagged as offensive language.
We present the offensive language detection results in the Figure  4, where the blue are safe and red are potentially offensive.
Figure 4. Distribution of offensive comments among language groups.
The dataset shows that offensive language is relatively rare, with only 0.43% of messages flagged as NSFW. This reflects typical user-generated platforms, where offensive content is present but remains a minority of overall interactions.
Hyperlinks
Links are naturally present in group conversations. However, these links could potentially contain unsafe content. Hence, we use regular expression (Thompson, 1968) detection to uncover 463,367 links in 313,906 messages and tag them for future modification.
3.3.2. Unsafe Text Replacement
  • Offensive Language Replacement: Offensive or harmful language detected in text messages is replaced with the placeholder label: detects this as offensive text. This approach preserves conversation structure while flagging offensive text, ensuring safety without losing context.
  • URL/Link Replacement: When URLs or links are identified in messages, they are replaced with the label: emoticonU detects a link. This ensures private or potentially harmful web addresses are not exposed in the dataset, while preserving the flow of conversation and preventing leakage of sensitive data.
3.4. Unsafe Image Moderation and Replacement
Next, we move on to unsafe image moderation, where we employ (1) detection and (2) text replacement.
3.4.1. Unsafe Image Detection
The purpose of image moderation is to remove offensive images; these include violent, sexual or offensive. To achieve this, we employ several readily available tools. We introduce them in the following;
  • nsfw-image-detection (AI, 2024): has over 54 million downloads on Huggingface and classifies images into normal and not safe for work (nsfw).
  • nsfw-classifier (Giacomo Arienti, 2024): classifies images into four categories - (1) drawings, (2) neutral, (3) hentai and (4) sexy.
  • vit-base-violence-detection (Haal, 2024): outputs binary labels for image violence detection.
  • vit-base-nsfw-detector (Codd, 2024): classifies images into nsfw and Safe For Work (sfw).
Figure 5. Distribution of nsfw vs. sfw amongst the four offensive image detection models on the raw data.
We utilize the frame processed above and static images into the four models, where we obtain the nsfw to sfw ratio. We present the results in Figure 5. nsfw-image-detection is 0.88%, nsfw-classifier is 0.93%, vit-base-violence-detection is 27.27%, vit-base-nsfw-detector is 0.47%. Then, 471,517 (92.06%) emoticons do not contain offensive taggings and and 40,675 (7.94%) images are considered offensive. Finally, there are a remaining 67 conversation groups.
3.4.2. Unsafe Image Removal and Replacement
For the identified offensive images, we replace the textual content with the placeholder detects this as offensive image. This helps maintain the integrity of the dataset while ensuring that no harmful visuals, such as explicit or violent content, are retained, thus reducing the risk of privacy violations or inappropriate representations. Then, we remove the emoticon entirely from our emoticon database.
3.5. Privatization
In addition to safety, the privacy of the data is paramount.
3.5.1. User Identifier
Hashing sender identifier
To protect the user’s privacy, we anonymize the user identifiers by hashing their original integer identifiers using the SHA-256 hash function (Appel, 2015). Originally, the user identifiers were integers ranging from five-digital value to a maximum value of ten-digit value. After applying the SHA-256 hashing algorithm (Appel, 2015), the user identifiers are transformed into 256-bit strings, which are irreversible and do not retain any direct correlation with the original numeric values. This ensures that the users’ identities are kept private, while still allowing for secure and effective data processing and analysis. one more
Replacing Mentions
It is common for users to interact with each other in the conversation group. To protect user privacy, we use regular expressions (Thompson, 1968) to identify user identifiers and mentions (e.g., username) within text and replace them with the label #USER. Additionally, when users are mentioned via hyperlinks, these occurrences are replaced with #USER_ID. Importantly, we do not replace the entire utterance but rather focus on replacing only the sensitive information, thus preserving the majority of the conversational context.
3.5.2. Message Information
We also anonymize message identifiers through the same SHA-256 hashing algorithm mentioned above.
3.5.3. Other Sensitive Information
To prevent the leakage of sensitive user information such as names, ages, addresses, organizations, or phone numbers, we use predefined context dictionaries to search the conversation logs. These logs are then further tagged by Llama-3.1-8B-Instruct (Touvron et al., 2023). Any sensitive information found is replaced with the label #SENSITIVE-INFORMATION within the conversation, but are not replaced entirely.
3.6. Manual Verification
After implementing the automatic process unsafe text, image and privatization we conduct a manual review as an additional safeguard to ensure the dataset’s safety and quality. We perform this in three phases (1) browsing through the conversation for private information leakage and, (2) relabeling them with appropriate tags.
3.7. Conversation Topic Labeling
To enable groups for future scenario analysis, we categorize conversations into domains. We utilize a two-step approach. First, we read the conversation name, which often provides a clear indication of the conversation’s general focus. Next, we browse through the conversation content, examining the topics and discussions that occur within. This process enables us to accurately identify the predominant subject matter of the conversation.
Our categorization is based on domains related to common hobbies and interests. We identify ten primary domains:
  • Language: This domain includes conversations focused on learning new languages, sharing language learning resources, or discussing language-related topics.
  • Arts: Conversations under this domain typically involve the sharing of artwork, discussions around different forms of art, drawing techniques, and other creative activities.
  • Games: Conversations in this domain revolve around various forms of gaming, whether they are video games, board games, or other interactive entertainment.
  • Technology: Conversations here engage in code sharing, learning new programming languages, discussing software development, or delving into the latest technological innovations.
  • Finance: This domain focuses on topics related to financial discussions, including traditional investing, cryptocurrency, blockchain technologies, and financial exchanges.
  • Social: Conversations under the Social domain are centered around meeting new people, making friends, or organizing social events such as meetups and hangouts.
  • Media Sharing: Conversations in this category often focus on sharing and discussing media content such as emoticons, memes, gifs, and videos for entertainment and social interaction.
  • Outdoor: Conversations here revolve around nature, outdoor activities like hiking and camping, and the appreciation of animals and the environment.
  • Anime: This domain is dedicated to discussions about anime, cartoons, and related pop culture topics, including anime recommendations and fan theories.
  • Fan Club: Conversations in this category are focused on specific celebrities, idols, or fictional characters, where users come together to share their admiration and discuss fan-related topics.
While these categories cover a wide range of interests, some topics naturally overlap. For instance, the increasingly popular genre of crypto-based gaming involves elements of both gaming and finance, as users play games to earn cryptocurrency. However, to maintain a concise and organized dataset, we prioritize the dominant domain in such cases—crypto gaming would thus fall under the Finance category in this example. The final dataset, after domain labeling, is presented and analyzed in the following section, where we examine the distribution of conversations across these domains and the insights derived from this classification.
4. Dataset Analysis
We analyze our dataset and share our findings in the following section. Firstly, we present a statistical overview of the dataset. Secondly, we dive into the multi-domain characteristics of our dataset. Finally, we analyze potential user behavior found in our dataset.
4.1. Statistic
4.1.1. Overall Information
Table 2. Statistics of the raw data and final emoticon dataset. Ours is the largest in size and contains user IDs across conversations.
Field Raw Data Ours
# Domains - 10
# Groups 70 67
# Unique Emoticons 142,981 105,933
Messages
# emoticons 512,192 370,222
# Text 8,332,351 8,286,422
# Avg. Emoticons per User 21.01 16.90
As shown in Table 2, our dataset is large in size and contains user information for cross conversation and cross domain analysis and modeling. Notably, we filter the raw data from the public platform for safety and privacy concerns to form the final dataset.
Table 3. Domain Statistics.
Category Text Messages Emoticons Users Sparsity Dom. Language
Global Unique Avg. Per User Global Use Emoticons
Anime 364,543 87,833 22,572 44.90 10,448 1,956 0.001989 French
Arts 19,515 219 176 3.37 328 65 0.019143 English
Fan Club 551,409 2,287 1,234 13.61 6,341 168 0.011032 Turkish
Finance 2,013,689 45,655 10,083 5.78 245,020 7,894 0.000574 Chinese
Game 2,289,100 79,367 16,382 32.16 48,297 2,468 0.001963 English
Language 2,388,938 51,556 15,801 14.92 50,734 3,456 0.000944 German
Outdoor 353,789 2,088 953 7.88 6,230 265 0.008268 English
Social 416,469 13,077 3,894 11.12 42,180 1,176 0.002856 English
Media Sharing 176,581 87,827 45,451 11.71 14,309 7,498 0.000258 English
Technology 20,305 313 210 4.17 4,425 75 0.019873 English
4.1.2. Emoticon Usage Distribution
Figure 6. Our Dataset Emoticon Usage Distribution. Number of users versus the number of emoticons. Our dataset emoticon usage follows a tail-shaped distribution.
Next, we analyze the distribution of emoticon usages. To better visualize this distribution and prevent distortion, we apply a logarithmic transformation to the number of users. As demonstrated in Figure  6, Our dataset exhibits a tail-shaped distribution, highlighting the real-world emoticon usage patterns.
4.1.3. Language Distribution
Then, we analyze the language distribution of our dataset, shown in Figure  7. Our dataset features Arabic, Bulgarian, Chinese, Dutch, English, French, German, Hindi, Italian, Japanese, Modern Greek, Polish, Portuguese, Russian, Spanish, Swahili, Thai, Turkish, Urdu, and Vietnamese. English is the most dominant language with an domineering 62.6%. Russian (9.6%) is the second most popular language followed by French (8.6%) and Chinese (8.0%).
Figure 7. Dataset Language Distribution. Our dataset features 20 languages.
4.2. Multi-Topic Characteristics
In this section, we analyze the multi-domain characteristics of our dataset. Firstly, we analyze the domain-group distribution. Secondly, we introduce the domain statistics. Lastly, we analyze the user-emoticon usage overlap across domains.
4.2.1. Group Distribution
As mentioned above, our dataset contains 67 groups and 10 domains, with the median group being five. We illustrate this in Figure 8. The minimum group count is one, being Arts, and the maximum group count is media sharing. This is due to the nature of resource, focusing on emoticons. Games and Social come close as they tend to be relaxed and so more expressed.
Figure 8. Our dataset features 67 groups and 10 domains. The median group count per domain is five.
4.2.2. Domain Statistics
Table 3 shows the domain overview statistics. Anime has the most emoticons, followed by Media Sharing and Game, reflecting the expressive nature of these groups. Technology is the sparsest domain, while Media Sharing is the densest, with Finance and Language being unexpectedly dense as well. The average emoticons per user are highest in Anime (44.90) and Game (32.16), while Technology (4.17) and Finance (5.78) have the lowest, suggesting that the nature of the domain influences emoticon usage.
4.2.3. User Emoticon Usage Overlap Across Domains
Lastly, we analyze the appearance of users in the same domain but in other groups, or in entirely different domain. We plot our findings in Figure 9.
Figure 9. Distribution of users’ emoticon usage.
Firstly, we establish that users do join other groups and participate; these could be within the same domain or an entirely different domain. The number beside the label on the horizontal axis reveals the total number of emoticon users within the domain, the blue bar represents the percentage of users that appear in other groups within the same domain and the red represents the percentage of users that appear in domains entirely different. We omit domains with fewer than 50 user emoticons to avoid generalization.
Secondly, two of the largest groups with more than 7000 emoticon users (Finance and Media Sharing domain) with roughly 15% and 3% to 85% and 97% different to same domain ratios. In fact, other than the Social domain, domains all have more users participating in the same domain than somewhere else; which could possibly reveal user preference and taste. Naturally, Social domain is clearly an explorative domain and therefore behaves as such.
5. Application Case 1: User Behavior Analysis
Next, we move on to the user behavior analysis on the dataset. More specifically, we demonstrate our dataset’s potential on showcasing (1) user distinct style (2) multi-domain behavior (3) temporal behavior change (4) user behavior with different respondents.
5.1. Distinct User Styles
Figure 10. User A emoticons in the Language group.
Figure 11. User B emoticons in the Game group.
As shown in Figure 10, User A’s emoticon styles are predominantly cat-themed and cute, while User B’s emoticons are more idol-based, with an emphasis on adorable and charming designs, as shown in Figure 11. This demonstrates that users generally have a distinct emoticon preference and this potentially be used for user modeling.
5.2. Cross Domain Behavior Change
More interestingly, we analyze if user, due to different context, could change their style. Herein, the same User A, Figure 12, displays a significantly different style in the Media Sharing group, which is more structured and cartoonish in appearance. This is contrastive to the originally cartoonish style.
Figure 12. User A emoticons in the Media Sharing group.
Figure 13. User B emoticons in the Outdoor group.
On the other hand, User B’s emoticons, Figure 13, still retain elements of their signature adorable style but are subtly tuned to match the outdoor-themed context of the new domain, which is about rats. This behavior reveals that different user have different consistency level in emoticon-usage. There are possibly many factors for this behavior such as the group context.
5.3. Temporal Behavior Change
Figure 14. User C emoticon usage over time where count is more than three.
We collate the top emoticon by usage count from User C and present in Figure  14. As can be seen, the user’s emoticon preferences evolve over time. In September and October, the user gravitates toward dog-themed, cartoonish emoticons. However, by January, their preference shifts to smaller, cuter dog emoticons. Interestingly, in June, the user shows a renewed interest in a different set of realistic dog emoticons, while still maintaining a fondness for the smaller, cuter dog designs.
Overall, despite fluctuations in style, the user consistently displays a preference for dog emoticons, with a notable emphasis on the cute, cartoonish variations. This suggests that while the user’s style may vary, their core preference remains for dog-themed emoticons, particularly those with a cute and cartoonish aesthetic.
Figure 15. User D emoticon usage over time. The user uses the green frog emoticon frequently overtime.
On the other hand, User D demonstrates a consistent use of a certain green laughing frog, except the months where no emoticons are used, as seen in Figure  15. Herein, the user demonstrates a consistent usage of a certain emoticon, unlike User C.
5.4. User Behavior with Different Respondents
Figure 16. User E shows a consistent style preference when replying to Users F and G.
We present two distinct cases of user behavior in response to different users. In the first case, User E replies to both Users F and G using a similar set of emoticons. As shown in Figure  16, User E consistently selects a boy emoticon with a nonchalant expression, regardless of the recipient. Although the third and fourth emoticons in the sequence differ, the first two emoticons remain frequently used, with counts of 12 and 8 respectively, indicating a strong preference for these emoticons.
Figure 17. User H shows contrasting style preferences when replying to Users I and J.
In contrast, User H exhibits a clear difference in emoticon usage when replying to Users I and J. As illustrated in Figure  17, User H uses bubbly, cheerful, and generally cute emoticons when replying to User I. However, when replying to User J, User H switches to a more melancholic and indifferent style, favoring stickman emoticons that are notably less cute. This shift in emoticon preference highlights the variation in user behavior based on the recipient of the message.
In conclusion, these examples demonstrate that the dataset captures meaningful differences in emoticon usage across various contexts. This variation, influenced by the person being replied to, suggests the potential for deeper personalization and more insightful user behavior analysis.
6. Application Case 2: Emoticon Recommendation
In this section, we evaluate the feasibility of our dataset through quantitative experiments. Specifically, we evaluate our dataset on current emoticon recommendation.
6.1. Experiment Settings
We evaluate two baseline methods for emoticon recommendation, described as follows:
  1. (1)
    MOD (Fei et al., 2021): Takes a list of negative samples along with one true positive sample and calculates the probability of each sample being the correct match. The results are then used to rank the samples.
  2. (2)
    SRS (Gao et al., 2021): This approach ranks a list containing negative and one positive sample, returning the ranked results.
We conduct experiments on three distinct perspectives of the dataset: (1) the English and Chinese subsets, (2) the domain-specific subset, and (3) the complete dataset.
6.1.1. English and Chinese Subset
Since MOD can only handle English and Chinese, we select groups with more than 80% of text being either English or Chinese and train the model on it. We do the same for SRS.
6.1.2. Domain
To interpretate the performance between different domains, we separate the dataset by domains as shown in Table 3. Since SRS is not limited by multilingualism, we evaluate using SRS.
6.1.3. Complete
Finally, the entire dataset is utilized for evaluation on SRS.
6.2. Results
6.2.1. English and Chinese Subset
Based on the experimental results presented in Table 4, we observe that the English language performs better than the Chinese language across both the MOD and SRS methods. Specifically, the English subset consistently shows higher scores in all evaluation metrics (MRR, R@1, R@3, R@5, and R@10). For instance, in the 1:9 ratio, the MOD method for English achieves an MRR of 0.289, while the Chinese MOD method only reaches 0.317, this could be due to the smaller data size of Chinese in comparison to English.
Table 4. emoticon Retrieval Results for MOD and SRS (English and Chinese). We shortform Negative Samples for NS.
Lang NS Method MRR R@1 R@3 R@5 R@10
En 1:9 MOD 0.289 0.307 0.495 0.224 1.000
SRS 0.291 0.098 0.298 0.495 1.000
1:19 MOD 0.180 0.051 0.145 0.250 0.511
SRS 0.299 0.050 0.155 0.256 0.429
Zh 1:9 MOD 0.317 0.125 0.344 0.531 1.000
SRS 0.264 0.056 0.281 0.500 1.000
1:19 MOD 0.156 0.063 0.094 0.156 0.406
SRS 0.270 0.061 0.138 0.219 0.377
6.2.2. Domain
Table 5 reveal that the emoticons domain performs the best overall, with an MRR of 0.335, R@1 of 0.143, R@3 of 0.365, and R@5 of 0.552. On the other hand Arts domain has a relatively lower performance, possibly due to the higher sparsity (0.019143).
Table 5. Results of SRS on individual domains.
Domain MRR R@1 R@3 R@5
Anime 0.294 0.101 0.298 0.500
Arts 0.288 0.094 0.281 0.484
Fan Club 0.280 0.080 0.280 0.510
Finance 0.296 0.080 0.280 0.510
Game 0.293 0.100 0.301 0.498
Language 0.296 0.103 0.307 0.501
Outdoor 0.298 0.103 0.324 0.500
Social 0.300 0.105 0.305 0.513
Media Sharing 0.335 0.143 0.365 0.552
Tech 0.321 0.125 0.361 0.611
6.2.3. Complete Dataset
The results in Table 6 show that performance decreases as the number of negative samples increases. With 4 negative samples, the model achieves a MRR of 0.425 and R@1 of 0.202, indicating good performance. As the number of negative sample increase, a decline in accuracy is observed, which is consistent with general recommendation models.
Table 6. Performance metrics for varying negative samples across multiple evaluation criteria on the entire dataset using SRS method.
NS Evaluation Metrics
MRR R@1 R@3 R@5 R@10 R@20
1:4 0.425 0.202 0.603 1.0 - -
1:9 0.294 0.102 0.303 0.499 1.0 -
1:19 0.296 0.054 0.153 0.251 0.332 1.0
7. Discussions
7.1. Other Potential Applications
Beyond user behavior analysis and emoticon recommendation, the dataset presents several opportunities for further research. One potential avenue is a more extensive User Behavior Modeling, which could involve analyzing temporal changes in user behavior and investigating whether emoticon usage varies based on the responder. Additional potential applications include, but are not limited to, Personalized Emoticon Retrieval and Generative AI, both of which have previously faced dataset limitations (Shen et al., 2024; Xu et al., 2024).
We acknowledge that there are several limitations in our work. Despite extensive automatic and manual data verification, privacy and safety concerns may still arise, presenting an opportunity for advancements in addressing such issues.
7.2. Ethical Considerations
We affirm that we have used the data in accordance with Telegram’s Terms and Conditions (tel, [n. d.]). Given a user’s expression of opposition, we refrain from crawling to respect their privacy. Additionally, our dataset aligns with the FAIR principles, which emphasize making data Findable, Accessible, Interoperable, and Reusable (Wilkinson et al., 2016).
8. Conclusion
In this paper, we introduce our emoticon dataset, the largest emoticon dataset to date, comprising data from 22K users, 370.2K emoticons, and 8.3M conversation messages. This multi-domain dataset encompasses a wealth of diverse information, capturing temporal, multilingual, and cross-domain behaviors not found in previous datasets. Extensive quantitative and qualitative experiments demonstrate the practical applications of our dataset in user behavior modeling and personalized emoticon recommendation. It also holds potential for further research in areas such as personalized retrieval and conversational studies.

### v2

@@
-A 106K Multi-Topic Multilingual Conversational User Dataset with Emoticons
+U-Sticker: A Large-Scale Multi-Domain User Sticker Dataset for Retrieval and Personalization
 Abstract.
-Instant messaging has become a predominant form of communication, with texts and emoticons enabling users to express emotions and ideas efficiently.
-Emoticons, in particular, have gained significant traction as a medium for conveying sentiments and information, leading to the growing importance of emoticon retrieval and recommendation systems.
-However, one of the key challenges in this area has been the absence of datasets that capture both the temporal dynamics and user-specific interactions with emoticons, limiting the progress of personalized user modeling and recommendation approaches.
-To address this, we introduce the emoticon dataset, a comprehensive resource that includes time-based data along with anonymous user identifiers across different conversations.
-As the largest publicly accessible emoticon dataset to date, it comprises 22K unique users, 370K emoticons, and 8.3M messages.
-The data was collected from a widely-used messaging platform across 67 conversations and 720 hours of crawling.
-Strict privacy and safety checks were applied to ensure the integrity of both text and image data.
-Spanning across 10 distinct domains, the emoticon dataset provides rich insights into temporal, multilingual, and cross-domain behaviors, which were previously unavailable in other emoticon-based datasets.
-Our in-depth experiments—both quantitative and qualitative—demonstrate the dataset’s potential in modeling user behavior and personalized recommendation systems, opening up new possibilities for research in personalized retrieval and conversational AI.
-The dataset is freely accessible  [^1] https://huggingface.co/datasets/metchee/u-sticker.
+Instant messaging with texts and stickers has become a widely adopted communication medium, enabling efficient expression of user semantics and emotions.
+With the increased use of stickers conveying information and feelings, sticker retrieval and recommendation has emerged as an important area of research.
+However, a major limitation in existing literature has been the lack of datasets capturing temporal and user-specific sticker interactions, which has hindered further progress in user modeling and sticker personalization.
+To address this, we introduce User-Sticker, a dataset that includes temporal and user anonymous ID across conversations.
+It is the largest publicly available sticker dataset to date, containing 22K unique users, 370K stickers, and 8.3M messages.
+The raw data was collected from a popular messaging platform from 67 conversations over 720 hours of crawling.
+All text and image data were carefully vetted for safety and privacy checks and modifications.
+Spanning 10 domains, the U-Sticker dataset captures rich temporal, multilingual, and cross-domain behaviors not previously available in other datasets.
+Extensive quantitative and qualitative experiments demonstrate U-Sticker’s practical applications in user behavior modeling and personalized recommendation and highlight its potential to further research areas in personalized retrieval and conversational studies.
+U-Sticker dataset is publicly available [^1] https://huggingface.co/datasets/metchee/u-sticker.
 Resource, Sticker Retrieval, User Modeling
-copyright: noneconference: Preprint; 2025; $c$s: Information systems Information retrievalccs: Information systems Dataset Personalization
+journalyear: 2025conference: the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13-18, 2025; Padua, Italyjournalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730311isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Information retrievalccs: Information systems Dataset Personalization
 1. Introduction
-Figure 1. The colorful and vivid emoticons makes for an expressive and comfortable conversation.
 Instant messaging (IM) has become an essential mode of communication, allowing users to interact efficiently.
-Beyond text, colorful and animated emoticons have emerged as a powerful medium of expression, enabling users to convey semantics and emotions comfortably and accurately (Ge, 2020; Constantin et al., 2019; Kariko and Anasih, 2019; Gygli and Soleymani, 2016; Jiang et al., 2023), as seen in Figure 1.
-Unlike general image datasets, which emphasize object recognition, emoticons carry rich semantic meaning, expressiveness, and emotional nuance, making their retrieval and recommendation fundamentally different from traditional image retrieval and recommendation tasks.
-However, despite emoticons’ widespread use, personalized emoticon retrieval and recommendation remains an understudied area, primarily due to the lack of large-scale user emoticon interaction dataset.
-As shown in Table 1, while several emoticon dataset exist, they often lack user information, making them unable to identify the same user across different conversations.
-Additionally, they contain few emoticons per conversation, providing limited insight into user preferences or emoticon usage patterns.
-Furthermore, some datasets are not open-sourced, limiting reproducibility and progress of personalized emoticon retrieval and recommendation.
+Beyond text, stickers have emerged as a powerful medium of expression, enabling users to convey semantics and emotions comfortably and accurately (Zhao et al., 2023; Fei et al., 2021).
+Unlike general image datasets, which emphasize object recognition, stickers carry rich semantic meaning, expressiveness, and emotional nuance, making their retrieval and recommendation fundamentally different from traditional image retrieval and recommendation tasks.
+However, despite stickers’ widespread use, personalized sticker retrieval and recommendation remains an understudied area, primarily due to the lack of large-scale user sticker interaction dataset.
+As shown in Table 1, while several sticker dataset exist, they often lack user information, making them unable to identify the same user across different conversations.
+Additionally, they contain few stickers per conversation, providing limited insight into user preferences or sticker usage patterns.
+Furthermore, some datasets are not open-sourced, limiting reproducibility and progress of personalized sticker retrieval and recommendation.
 Table 1.
-Comparison with Emoticon Datasets.
+Comparison with Stickers Datasets.
 We bold the largest count among publicly available datasets and underline the second largest.
-Our dataset is cross-domain, large in scale and contain time and user ID for personalization.
-Dataset Pub. Avail. Cross Domain Time Info. # Sticks./user # Emoticons Description
-StickerTag (Wang et al., 2024) 13,571 Emoticon-tag pairs
-StickerInt (Liang et al., 2024) Unknown 1,025 Dialogues with emoticons
-StickerCLIP (Zhao et al., 2023) 820,000 Chinese emoticon-tag pairs
-CSMSA (Ge et al., 2022) 16,000 Emoticons-tag pairs
-SRS, PESRS (Gao et al., 2021, 2020) 6.82 320,168 Dialogues with emoticons
-PerSRV (Chee et al., 2024) Unknown 543,098 Emoticon-query pairs
-MCDSCS (Shi and Kong, 2024) 2.03 14,400 Dialogues with emoticons
-SER30K (Liu et al., 2022) 30,739 Emoticon-sentiments pairs
-MOD (Fei et al., 2021) 1.36 307 Dialogues with emoticons
-Ours 16.90 370,222 Dialogues with emoticons
-To address these limitations, we introduce a new emoticon dataset, the first large-scale dataset that includes both user information and emoticon-based conversations.
-Our contributions can be summarized as follows:
-  • We present a new emoticon dataset, the largest emoticon dataset to date, containing 22k users, 370.2k emoticons and 8.3M conversation messages of texts and emoticons.
-  • Our dataset is a multi-domain dataset that includes rich and diverse information.
+Our dataset is cross-domain, large in scale, and contains time and user ID for personalization.
+Dataset Pub. Avail. Cross Domain Time Info. # Sticks./user # Stickers # Languages
+StickerTag (Wang et al., 2024) 13,571 1
+StickerInt (Liang et al., 2024) Unknown 1,025 1
+StickerCLIP (Zhao et al., 2023) 820,000 1
+CSMSA (Ge et al., 2022) 16,000 1
+SRS, PESRS (Gao et al., 2021, 2020) 6.82 320,168 1
+PerSRV (Chee et al., 2024) Unknown 543,098 1
+MCDSCS (Shi and Kong, 2024) 2.03 14,400 2
+SER30K (Liu et al., 2022) 30,739 1
+MOD (Fei et al., 2021) 1.36 307 2
+Ours 16.90 370,222 18
+To address these limitations, we introduce U-Sticker, the first large-scale dataset that includes both user information and sticker-based conversations. Our contributions are summarized as follows:
+  • We present U-Sticker, the largest sticker dataset to date, containing 22K users, 370K stickers and 8.3M conversation messages of texts and stickers.
+  • U-Sticker is a multi-domain dataset that includes rich and diverse information.
 Covering 10 domains, it captures temporal, multilingual, and cross-domain behaviors that are not present in previous datasets.
-  • Extensive quantitative and qualitative experiments demonstrate the dataset’s practical applications on user behavior analysis and modeling, personalized emoticon recommendation.
+  • Extensive quantitative and qualitative experiments demonstrate U-Sticker’s practical applications on user behavior analysis and modeling, personalized sticker recommendation.
 It also holds potential for further research in areas such as personalized retrieval and conversational studies.
 2. Related Work
-We outline the current emoticon datasets and their applications.
-The terms sticker, emoticon, memes, gif mean approximately the same, dynamic or static multimedia that can be shared to express emotions.
-Herein, we use the word emoticon.
-2.1. Emoticon Datasets
-Due to the challenges in preparing emoticon conversation-based datasets, their availability remains limited, particularly for publicly accessible datasets.
-However, recent research on emoticons has gained more attention.
+2.1. Sticker Datasets
+Due to the challenges in preparing sticker conversation-based datasets, their availability remains limited, particularly for publicly accessible datasets.
+However, recent research on stickers has gained more attention.
 Currently, there are only nine related datasets, with only three being publicly available.
@@
 Specifically, six datasets are not publicly accessible (Wang et al., 2024; Liang et al., 2024; Zhao et al., 2023; Ge et al., 2022; Gao et al., 2021; Chee et al., 2024), meaning they are either closed-source or require additional access that is often unattainable.
-This lack of open access presents a significant challenge for emoticon-based research.
-2.1.2. Multi Topic Information
+This lack of open access presents a significant challenge for sticker-based research.
+2.1.2. Cross Domain Information
 Furthermore, none of the aforementioned datasets contain cross-domain information, as the conversations or dialogues are not identified with specific topics.
-This limitation hinders the exploration of cross-domain emoticon usage, which could provide valuable insights.
+This limitation hinders the exploration of cross-domain sticker usage, which could provide valuable insights.
 2.1.3. User Information
-Additionally, publicly available datasets are scarce in terms of user data, with 1.36 (Fei et al., 2021) and 2.03 (Shi and Kong, 2024) historical emoticons per user.
+Additionally, publicly available datasets are scarce in terms of user data, with 1.36 (Fei et al., 2021) and 2.03 (Shi and Kong, 2024) historical stickers per user.
 This lack of data makes it difficult to perform user analysis and personalization-related tasks.
-Our dataset bridges this gap, providing the highest emoticon history per user, with an impressive 16.90 emoticons.
+U-Sticker bridges this gap, providing the highest sticker history per user, with an impressive 16.90 stickers.
 2.1.4. Multilingualism
-Most existing datasets focus on one or two languages, limiting the ability of smaller communities to benefit from advancements in emoticon personalization.
-In contrast, our dataset is multilingual, allowing a broader range of communities to engage with these advancements.
-As shown in Table 1 and our analysis, our dataset is the largest publicly available emoticon-user dataset.
-It features 370,222 emoticons, cross-domain behavior, rich contextual information, and comprehensive user and temporal data.
-2.2. Emoticon Retrieval
-Most previous research emphasizes the importance of data for emoticon retrieval.
+Most existing datasets focus on one or two languages, limiting the ability of smaller communities to benefit from advancements in sticker personalization.
+In contrast, U-Sticker is multilingual, allowing a broader range of communities to engage with these advancements.
+As see in Table 1 and above, U-Sticker is the largest available sticker-user dataset.
+It features 370,222 stickers, cross-domain behavior, rich contextual information, and comprehensive user and temporal data.
+2.2. Sticker Retrieval
+Most previous research emphasizes the importance of data for sticker retrieval.
 SRS (Gao et al., 2020), PESRS (Gao et al., 2021) require corresponding utterances, while Lao et al.
 (Lao et al., 2022) rely on manually labeled emotions, sentiments, and reply keywords.
-CKES (Chen et al., 2024) annotates each emoticon with a corresponding emotion.
-During emoticon creation, Hike Messager (Laddha et al., 2020) tags conversational phrases to emoticons.
-The reliance on data presents a significant limitation, as emoticons without associated information are excluded from consideration.
+CKES (Chen et al., 2024)annotates each sticker with a corresponding emotion.
+During sticker creation, Hike Messager (Laddha et al., 2020) tags conversational phrases to stickers.
+The reliance on data presents a significant limitation, as stickers without associated information are excluded from consideration.
 Gao et al.
-(Gao et al., 2020) use a convolutional emoticon encoder and self-attention dialog encoder for emoticon-utterance representations, followed by a deep interaction network and fusion network to capture dependencies and output the final matching score.
-The method selects the ground truth emoticon from a pool of emoticon candidates and its successor.
+(Gao et al., 2020) use a convolutional sticker encoder and self-attention dialog encoder for sticker-utterance representations, followed by a deep interaction network and fusion network to capture dependencies and output the final matching score.
+The method selects the ground truth sticker from a pool of sticker candidates and its successor.
 Zhang et al.
@@
 The PBR (Xia et al., [n.
-d.]) paradigm enhances emotion comprehension through knowledge distillation, contrastive learning, and improved hard negative sampling to generate diverse and discriminative emoticon representations for better response matching in dialogues.
-PEGS (Zhang et al., 2024), StickerInt (Liang et al., 2024) generate emoticon information using multimodal models and selects emoticon responses, but does not consider personalization.
+d.]) paradigm enhances emotion comprehension through knowledge distillation, contrastive learning, and improved hard negative sampling to generate diverse and discriminative sticker representations for better response matching in dialogues.
+PEGS (Zhang et al., 2024), StickerInt (Liang et al., 2024) generate sticker information using multimodal models and selects sticker responses, but does not consider personalization.
 StickerCLIP (Zhao et al., 2023) fine-tunes pretrained image encoders but does not consider personalization.
-In addition, many methods designed to rank emoticons from top-k candidates face a significant drawback in real-world emoticon retrieval scenarios, as they quickly become impractical when applied to larger datasets.
+In addition, many methods designed to rank stickers from top-k candidates face a significant drawback in real-world sticker retrieval scenarios, as they quickly become impractical when applied to larger datasets.
 Possibly due to the lack of dataset, most methods do not consider personalization.
-2.3. Personalized Emoticon Recommendation
-The absence of personalized emoticon recommendation can likely be attributed to the scarcity of emoticon-user datasets.
+2.3. Personalized Sticker Recommendation
+The absence of personalized sticker recommendation can likely be attributed to the scarcity of sticker-user datasets.
 PESRS (Gao et al., 2021) improves upon earlier work (Gao et al., 2020) by incorporating user preferences, but the dataset it relies on is not publicly accessible.
-This limited availability indirectly hampers the development of personalized emoticon recommendations and may impede further advancements in the field.
-3. Our Dataset
-Telegram (tel, [n.
-d.]) is a widely used open-source messaging platform, offering publicly accessible conversations.
+This limited availability indirectly hampers the development of personalized sticker recommendations and may impede further advancements in the field.
+3. U-Sticker Dataset
+Telegram is a widely used open-source messaging platform, offering publicly accessible conversations.
 The interactions within these groups are natural, making them a valuable source for understanding human behavior, particularly in the context of personalization.
-As such, they present a promising dataset for emoticon-related research (Dargahi Nobari et al., 2017; Naseri and Zamani, 2019; Baumgartner et al., 2020; Blas et al., 2024; Khaund et al., 2020; La Morgia et al., 2025).
+As such, they present a promising dataset for sticker-related research.
 However, the raw data requires preprocessing before it can be utilized effectively.
@@
 Given the significant implications of this dataset for future research, we adhere to the following criteria to ensure the quality and reliability of the data source
-We firstly define the envision for the final user-emoticon dataset.
-Since emoticons are the focal point of this dataset, we would like a high presence of emoticons.
+We firstly define the envision for the final user-sticker dataset.
+Since stickers are the focal point of this dataset, we would like a high presence of stickers.
 Additionally, since we would like to analyze interactions across different domains, the final dataset should involve diverse topics.
@@
   1. (1)
-    Emoticon prevalence: Conversations must contain a significant number of emoticons.
+    Sticker prevalence: Conversations must contain a significant number of stickers.
   2. (2)
@@
     Scalability: We aim to capture as many suitable conversations as possible.
-Following these criteria, we manually screened hundreds of conversation groups and ultimately selected 71 conversation groups to crawl their content.
+Following these criteria, we manually screened hundreds of conversation groups and ultimately selected 70 conversation groups to crawl their content.
 3.2. Dataset Pre-processing
-The dataset, crawled using Telethon (Exo, 2025), contains key components from user messages, including the unique conversation group ID, group name, message datetime, user ID, media types, message ID, message text, and emoticon details (ID, set ID, access hash, and mime type).
-Users may participate in multiple conversation groups.
+We crawl the data using Telethon (Exo, 2025) and process the data in the following sections. Note that users can participate in multiple conversation groups.
 3.2.1. Text Processing
@@
 After language detection, we manually review the classification results, removing entries with fewer than 20 occurrences and those from administrative bots.
-The final language distribution is illustrated in Figure  2.
 In total, we detect 18 languages, with the top ten languages in order being English, Russian, French, Spanish, Chinese, Polish, Italian and German, Turkish, Portugese.
-As can be seen, emoticonU features 18 languages which is potentially viable for multilingualism emoticon-related task.
-Figure 2.
-Distribution of detected languages in the dataset.
-The dataset features over 18 languages.
-The top ten languages are English, Russian, French, Spanish, Chinese, Polish, Italian and German, Turkish, Portugese.
-3.2.2. Emoticon Processing
-The downloaded emoticons come in three formats; .webp, .webm, and .tgs, we present the extension distribution in Figure 3.
+As can be seen, U-Sticker features 18 languages which is potentially viable for multilingualism sticker-related task.
+3.2.2. Sticker Processing
+The downloaded stickers come in three formats; .webp, .webm, and .tgs, the distribution is 65.6%, 23.0% and 11.4% respectively.
 We convert all tgs files to the .gif extension using tgsconverter (Benau, 2025), .webp is converted to .png and no conversion is done to .webm.
 We break up dymamic images into frames and save a frame every second to prepare for later use.
-Figure 3. Preliminary emoticons extension distribution. 65.6% of the emoticons are static and 34.4% are dyanimc.
 3.3. Unsafe Text Detection and Replacement
@@
 For each, we employ well-established moderation models that have been published in peer-reviewed conferences, we describe the models in the following and the threshold values.
+The threshold values were manually selected through stepwise examination of the model outputs, aiming to balance the inclusion of high-quality text with the exclusion of plausibly offensive content.
   • roberta-base-cold (Deng et al., 2022): Detects offensive language in Chinese. The model returns a binary label, positive values are tagged as offensive language.
@@
   • dehatebert-mono-polish (Aluru et al., 2020): Detects hate speech in Polish language. The model returns a binary label, positive values are tagged as offensive language.
-We present the offensive language detection results in the Figure  4, where the blue are safe and red are potentially offensive.
-Figure 4. Distribution of offensive comments among language groups.
+We present the offensive language detection results in Figure  1.
+Figure 1. Offensive comment distribution across text classification models. 0.43% of comments are flagged as offensive.
 The dataset shows that offensive language is relatively rare, with only 0.43% of messages flagged as NSFW.
@@
 3.3.2. Unsafe Text Replacement
-  • Offensive Language Replacement: Offensive or harmful language detected in text messages is replaced with the placeholder label: detects this as offensive text.
+  • Offensive Language Replacement: Offensive or harmful language detected in text messages is replaced with the placeholder label: U-Sticker detects this as offensive text.
 This approach preserves conversation structure while flagging offensive text, ensuring safety without losing context.
-  • URL/Link Replacement: When URLs or links are identified in messages, they are replaced with the label: emoticonU detects a link.
+  • URL/Link Replacement: When URLs or links are identified in messages, they are replaced with the label: U-Sticker detects a link.
 This ensures private or potentially harmful web addresses are not exposed in the dataset, while preserving the flow of conversation and preventing leakage of sensitive data.
@@
   • vit-base-nsfw-detector (Codd, 2024): classifies images into nsfw and Safe For Work (sfw).
-Figure 5. Distribution of nsfw vs. sfw amongst the four offensive image detection models on the raw data.
+Figure 2.
+Distribution of NSFW vs.
+SFW amongst the four offensive image detection models on the raw data.
+To ensure final labale assignment robustness, we utilize a voting mechanism to balance individual model biases.
 We utilize the frame processed above and static images into the four models, where we obtain the nsfw to sfw ratio.
@@
 nsfw-image-detection is 0.88%, nsfw-classifier is 0.93%, vit-base-violence-detection is 27.27%, vit-base-nsfw-detector is 0.47%.
-Then, 471,517 (92.06%) emoticons do not contain offensive taggings and and 40,675 (7.94%) images are considered offensive.
-Finally, there are a remaining 67 conversation groups.
 3.4.2. Unsafe Image Removal and Replacement
-For the identified offensive images, we replace the textual content with the placeholder detects this as offensive image.
+For the identified offensive images, we replace the textual content with the placeholder U-Sticker detects this as offensive image.
 This helps maintain the integrity of the dataset while ensuring that no harmful visuals, such as explicit or violent content, are retained, thus reducing the risk of privacy violations or inappropriate representations.
-Then, we remove the emoticon entirely from our emoticon database.
+Then, we remove the sticker entirely from our sticker database.
 3.5. Privatization
@@
 This ensures that the users’ identities are kept private, while still allowing for secure and effective data processing and analysis.
-one more
 Replacing Mentions
@@
 To prevent the leakage of sensitive user information such as names, ages, addresses, organizations, or phone numbers, we use predefined context dictionaries to search the conversation logs.
-These logs are then further tagged by Llama-3.1-8B-Instruct (Touvron et al., 2023).
+These logs are then further tagged by Llama-1.5b (Liu et al., 2023).
 Any sensitive information found is replaced with the label #SENSITIVE-INFORMATION within the conversation, but are not replaced entirely.
@@
 After implementing the automatic process unsafe text, image and privatization we conduct a manual review as an additional safeguard to ensure the dataset’s safety and quality.
-We perform this in three phases (1) browsing through the conversation for private information leakage and, (2) relabeling them with appropriate tags.
-3.7. Conversation Topic Labeling
+In this phase, a subset of the dataset is randomly sampled and carefully examined by human reviewers.
+The goal is to verify that the automated procedures have successfully removed or replaced all sensitive and harmful content without affecting the overall structure or meaning of the messages.
+The manual review process is essential for catching any potential issues that automated processes may have missed, such as false negatives or inaccuracies in labeling.
+Finally, 370,222 (72.28%) of 512,192 stickers do not contain offensive taggings, 8,286,422 (99.45%) of 8,332,351 utterances are safe and private. Then, there are a remaining 67 conversation groups.
+3.7. Conversation Domain Labeling
 To enable groups for future scenario analysis, we categorize conversations into domains.
@@
   • Social: Conversations under the Social domain are centered around meeting new people, making friends, or organizing social events such as meetups and hangouts.
-  • Media Sharing: Conversations in this category often focus on sharing and discussing media content such as emoticons, memes, gifs, and videos for entertainment and social interaction.
+  • Media Sharing: Conversations in this category often focus on sharing and discussing media content such as stickers, memes, gifs, and videos for entertainment and social interaction.
   • Outdoor: Conversations here revolve around nature, outdoor activities like hiking and camping, and the appreciation of animals and the environment.
@@
 4. Dataset Analysis
-We analyze our dataset and share our findings in the following section.
+We analyze the U-Sticker dataset and share our findings in the following section.
 Firstly, we present a statistical overview of the dataset.
-Secondly, we dive into the multi-domain characteristics of our dataset.
-Finally, we analyze potential user behavior found in our dataset.
+Secondly, we dive into the multi-domain characteristics of U-Sticker.
+Finally, we analyze potential user behavior found in U-Sticker.
 4.1. Statistic
 4.1.1. Overall Information
-Table 2. Statistics of the raw data and final emoticon dataset. Ours is the largest in size and contains user IDs across conversations.
-Field Raw Data Ours
+Table 2. Statistics and Comparison of Raw Data and U-Sticker (A) and Top-8 Languages Post-Processing (B).
+A. Statistics Comparison.
+Field Raw Data U-Sticker
 # Domains - 10
 # Groups 70 67
-# Unique Emoticons 142,981 105,933
+# Unique Stickers 142,981 105,803
 Messages
-# emoticons 512,192 370,222
+# Stickers 512,192 370,222
 # Text 8,332,351 8,286,422
-# Avg. Emoticons per User 21.01 16.90
-As shown in Table 2, our dataset is large in size and contains user information for cross conversation and cross domain analysis and modeling.
-Notably, we filter the raw data from the public platform for safety and privacy concerns to form the final dataset.
-Table 3. Domain Statistics.
-Category Text Messages Emoticons Users Sparsity Dom. Language
-Global Unique Avg. Per User Global Use Emoticons
-Anime 364,543 87,833 22,572 44.90 10,448 1,956 0.001989 French
-Arts 19,515 219 176 3.37 328 65 0.019143 English
-Fan Club 551,409 2,287 1,234 13.61 6,341 168 0.011032 Turkish
-Finance 2,013,689 45,655 10,083 5.78 245,020 7,894 0.000574 Chinese
-Game 2,289,100 79,367 16,382 32.16 48,297 2,468 0.001963 English
-Language 2,388,938 51,556 15,801 14.92 50,734 3,456 0.000944 German
-Outdoor 353,789 2,088 953 7.88 6,230 265 0.008268 English
-Social 416,469 13,077 3,894 11.12 42,180 1,176 0.002856 English
-Media Sharing 176,581 87,827 45,451 11.71 14,309 7,498 0.000258 English
-Technology 20,305 313 210 4.17 4,425 75 0.019873 English
-4.1.2. Emoticon Usage Distribution
-Figure 6. Our Dataset Emoticon Usage Distribution. Number of users versus the number of emoticons. Our dataset emoticon usage follows a tail-shaped distribution.
-Next, we analyze the distribution of emoticon usages.
+Avg. Stickers/User 20.54 16.90
+B. Languages.
+Top-8 Languages
+1. English (62.6%)
+2. Russian (9.6%)
+3. French (8.6%)
+4. Chinese (8.0%)
+5. Spanish (2.5%)
+6. Turkish (2.0%)
+7. Vietnamese (1.9%)
+8. Polish (1.5%)
+As shown in Table 2, our U-Sticker is large in size and contains user information for cross conversation and cross domain analysis and modeling.
+Notably, we filter the raw data from the public platform for safety and privacy concerns to form the final U-Sticker dataset.
+Table 3.
+Domain Statistics.
+Notably, Anime, Media Sharing, and Game are the most active domains in terms of sticker usage, while Technology is the least active.
+The sparsity values highlight the varying levels of engagement across domain.
+Sparsity values are scaled by a factor of 10,000 for easier readability.
+Category Sticker Users Sparsity
+Global Unique Global Sticker
+Anime (4) 87,833 22,572 10,448 1,956 19.89
+Arts (1) 219 176 3.37 65 191.43
+Fan Club (4) 2,287 1,234 6,341 168 110.32
+Finance (8) 45,655 10,083 245,020 7,894 5.74
+Game (1) 79,367 16,382 48,297 2,468 19.63
+Language (5) 51,556 15,801 50,734 3,456 9.44
+Outdoor (5) 2,088 953 6,230 265 82.68
+Social (9) 13,077 3,894 42,180 1,176 28.56
+Media (18) 87,827 45,451 14,309 7,498 2.58
+Tech (2) 313 210 4,425 75 198.73
+4.1.2. Sticker Usage Distribution
+Figure 3. Sticker Usage Distribution. Number of users against the number of stickers. U-Sticker usage trend follows a long-tail distribution, highlighting the real-world sticker usage patterns.
+Next, we analyze the distribution of sticker usages.
 To better visualize this distribution and prevent distortion, we apply a logarithmic transformation to the number of users.
-As demonstrated in Figure  6, Our dataset exhibits a tail-shaped distribution, highlighting the real-world emoticon usage patterns.
+As demonstrated in Figure  3, the U-Sticker usage trend follows a long-tail distribution, where most users employ between 0 and 100 stickers, while only a small fraction exhibit extremely high usage.
+This behavior is consistent with Zipfian or Pareto-like distributions (Zipf, 1949) (new, [n.
+d.]), and is typical of organic user activity observed in large-scale, real-world datasets.s.
 4.1.3. Language Distribution
-Then, we analyze the language distribution of our dataset, shown in Figure  7.
-Our dataset features Arabic, Bulgarian, Chinese, Dutch, English, French, German, Hindi, Italian, Japanese, Modern Greek, Polish, Portuguese, Russian, Spanish, Swahili, Thai, Turkish, Urdu, and Vietnamese.
+Then, we analyze the language distribution of U-Sticker, shown in Figure  2.
+U-Sticker features Arabic, Bulgarian, Chinese, Dutch, English, French, German, Hindi, Italian, Japanese, Modern Greek, Polish, Portuguese, Russian, Spanish, Swahili, Thai, Turkish, Urdu, and Vietnamese.
 English is the most dominant language with an domineering 62.6%.
 Russian (9.6%) is the second most popular language followed by French (8.6%) and Chinese (8.0%).
-Figure 7. Dataset Language Distribution. Our dataset features 20 languages.
-4.2. Multi-Topic Characteristics
-In this section, we analyze the multi-domain characteristics of our dataset.
+4.2. Multi-Domain Characteristics
+In this section, we analyze the multi-domain characteristics of U-Sticker.
 Firstly, we analyze the domain-group distribution.
 Secondly, we introduce the domain statistics.
-Lastly, we analyze the user-emoticon usage overlap across domains.
+Lastly, we analyze the user-sticker usage overlap across domains.
 4.2.1. Group Distribution
-As mentioned above, our dataset contains 67 groups and 10 domains, with the median group being five.
-We illustrate this in Figure 8.
+As mentioned above, U-Sticker contains 67 groups and 10 domains, with the median group being five.
 The minimum group count is one, being Arts, and the maximum group count is media sharing.
-This is due to the nature of resource, focusing on emoticons.
+This is due to the nature of resource, focusing on stickers.
 Games and Social come close as they tend to be relaxed and so more expressed.
-Figure 8. Our dataset features 67 groups and 10 domains. The median group count per domain is five.
 4.2.2. Domain Statistics
-Table 3 shows the domain overview statistics.
-Anime has the most emoticons, followed by Media Sharing and Game, reflecting the expressive nature of these groups.
-Technology is the sparsest domain, while Media Sharing is the densest, with Finance and Language being unexpectedly dense as well.
-The average emoticons per user are highest in Anime (44.90) and Game (32.16), while Technology (4.17) and Finance (5.78) have the lowest, suggesting that the nature of the domain influences emoticon usage.
-4.2.3. User Emoticon Usage Overlap Across Domains
+We present the overview statistics of domain in Table 3.
+Looking at sticker global counts, the domain with the most sticker is Anime, followed by Media Sharing and Game.
+This is reasonable as these are highly relaxed groups and tend to be more expressive, thereby having a higher sticker usage count.
+The most sparse domain is Technology and the most dense domain is Media Sharing.
+This is natural as technology is a rather technical topic and audience could be less inclined to using sticker expressions.
+On the other hand, media sharing sometimes encourage chats entirely comprising of stickers, hence increasing the density.
+Surprisingly, Finance and Language are also denser domains.
+This could be because Finance and Language are semi-professional scenarios which could call for an increased sticker usage.
+Lastly, the average stickers used per user varies largely with Anime and then game domain having the highest 44.90 and 32.16 respectively.
+On the other hand, Technology and Finance domains have 4.17 and 5.78 stickers respectively.
+We hypothesize that the major difference in the nature of the scenario affects sticker usage in users.
+4.2.3. User Sticker Usage Overlap Across Domains
 Lastly, we analyze the appearance of users in the same domain but in other groups, or in entirely different domain. We plot our findings in Figure 4.
-Figure 9. Distribution of users’ emoticon usage.
+Figure 4. Distribution of users’ sticker usage. Social domain users features the largest distribution of users in contrastive (yellow), reflecting the inherently explorative nature of their domain.
 Firstly, we establish that users do join other groups and participate; these could be within the same domain or an entirely different domain.
-The number beside the label on the horizontal axis reveals the total number of emoticon users within the domain, the blue bar represents the percentage of users that appear in other groups within the same domain and the red represents the percentage of users that appear in domains entirely different.
-We omit domains with fewer than 50 user emoticons to avoid generalization.
-Secondly, two of the largest groups with more than 7000 emoticon users (Finance and Media Sharing domain) with roughly 15% and 3% to 85% and 97% different to same domain ratios.
+The number beside the label on the horizontal axis reveals the total number of sticker users within the domain, the blue bar represents the percentage of users that appear in other groups within the same domain.
+We omit domains with fewer than 50 user stickers to avoid generalization.
+Secondly, two of the largest groups with more than 7000 sticker users (Finance and Media Sharing domain) with roughly 15% and 3% to 85% and 97% different to same domain ratios.
 In fact, other than the Social domain, domains all have more users participating in the same domain than somewhere else; which could possibly reveal user preference and taste.
@@
 5. Application Case 1: User Behavior Analysis
-Next, we move on to the user behavior analysis on the dataset.
-More specifically, we demonstrate our dataset’s potential on showcasing (1) user distinct style (2) multi-domain behavior (3) temporal behavior change (4) user behavior with different respondents.
+Next, we move on to the user behavior analysis on the U-Sticker dataset. More specifically, we demonstrate U-Sticker’s potential on showcasing (1) user distinct style (2) multi-domain behavior.
 5.1. Distinct User Styles
-Figure 10. User A emoticons in the Language group.
-Figure 11. User B emoticons in the Game group.
-As shown in Figure 10, User A’s emoticon styles are predominantly cat-themed and cute, while User B’s emoticons are more idol-based, with an emphasis on adorable and charming designs, as shown in Figure 11.
-This demonstrates that users generally have a distinct emoticon preference and this potentially be used for user modeling.
+As shown in Figure 5, User A’s sticker styles are predominantly cat-themed and cute, while User B’s stickers are more idol-based, with an emphasis on adorable and charming designs.
+This demonstrates that users generally have a distinct sticker preference and this potentially be used for user modeling.
+Figure 5. User A stickers in the Language group and User B stickers in the Game group.
 5.2. Cross Domain Behavior Change
+Figure 6. User A stickers in the Media Sharing group and User B stickers in the Outdoor group.
 More interestingly, we analyze if user, due to different context, could change their style.
-Herein, the same User A, Figure 12, displays a significantly different style in the Media Sharing group, which is more structured and cartoonish in appearance.
-This is contrastive to the originally cartoonish style.
-Figure 12. User A emoticons in the Media Sharing group.
-Figure 13. User B emoticons in the Outdoor group.
-On the other hand, User B’s emoticons, Figure 13, still retain elements of their signature adorable style but are subtly tuned to match the outdoor-themed context of the new domain, which is about rats.
-This behavior reveals that different user have different consistency level in emoticon-usage.
+Herein, the same User A, Figure 6, displays a significantly different style in the Media Sharing group, which is more structured and cartoonish in appearance, vastly contrastive to the originally cartoonish style.
+On the other hand, User B’s stickers, still retain elements of their signature adorable style but are subtly tuned to match the outdoor-themed context of the new domain, which is about rats.
+This behavior reveals that different user have different consistency level in sticker-usage.
 There are possibly many factors for this behavior such as the group context.
 5.3. Temporal Behavior Change
-Figure 14. User C emoticon usage over time where count is more than three.
-We collate the top emoticon by usage count from User C and present in Figure  14.
-As can be seen, the user’s emoticon preferences evolve over time.
-In September and October, the user gravitates toward dog-themed, cartoonish emoticons.
-However, by January, their preference shifts to smaller, cuter dog emoticons.
-Interestingly, in June, the user shows a renewed interest in a different set of realistic dog emoticons, while still maintaining a fondness for the smaller, cuter dog designs.
-Overall, despite fluctuations in style, the user consistently displays a preference for dog emoticons, with a notable emphasis on the cute, cartoonish variations.
-This suggests that while the user’s style may vary, their core preference remains for dog-themed emoticons, particularly those with a cute and cartoonish aesthetic.
-Figure 15. User D emoticon usage over time. The user uses the green frog emoticon frequently overtime.
-On the other hand, User D demonstrates a consistent use of a certain green laughing frog, except the months where no emoticons are used, as seen in Figure  15.
-Herein, the user demonstrates a consistent usage of a certain emoticon, unlike User C.
-5.4. User Behavior with Different Respondents
-Figure 16. User E shows a consistent style preference when replying to Users F and G.
-We present two distinct cases of user behavior in response to different users.
-In the first case, User E replies to both Users F and G using a similar set of emoticons.
-As shown in Figure  16, User E consistently selects a boy emoticon with a nonchalant expression, regardless of the recipient.
-Although the third and fourth emoticons in the sequence differ, the first two emoticons remain frequently used, with counts of 12 and 8 respectively, indicating a strong preference for these emoticons.
-Figure 17. User H shows contrasting style preferences when replying to Users I and J.
-In contrast, User H exhibits a clear difference in emoticon usage when replying to Users I and J.
-As illustrated in Figure  17, User H uses bubbly, cheerful, and generally cute emoticons when replying to User I.
-However, when replying to User J, User H switches to a more melancholic and indifferent style, favoring stickman emoticons that are notably less cute.
-This shift in emoticon preference highlights the variation in user behavior based on the recipient of the message.
-In conclusion, these examples demonstrate that the dataset captures meaningful differences in emoticon usage across various contexts.
-This variation, influenced by the person being replied to, suggests the potential for deeper personalization and more insightful user behavior analysis.
-6. Application Case 2: Emoticon Recommendation
-In this section, we evaluate the feasibility of our dataset through quantitative experiments. Specifically, we evaluate our dataset on current emoticon recommendation.
+We present the top-used stickers of User C in Figure  7, showing how their preferences shift over time.
+In September and October, they favor cartoonish dog stickers; by January, they prefer smaller, cuter dog designs; and in June, they lean toward realistic dog stickers while still using the cute ones.
+Despite style changes, their consistent preference for dog-themed stickers—especially cute ones—remains.
+Figure 7.
+User C’s evolving sticker preferences over time, shifting from expressive dog stickers in September to smaller, more reserved dog stickers by January, with a renewed interest in realistic dog stickers by June, maintaining a consistent preference for dog-related designs.
+In contrast, User D, Figure  8 consistently uses the same green laughing frog sticker (commonly known as ”Pepe”) across many months, with few gaps.
+Overall, this demonstrates that our dataset contains both types of users: those like User D, who exhibit stable preferences for specific stickers, and those like User C, whose preferences evolve over time.
+Intuitively, this aligns with common sense, as people can have both consistent and fluctuating preferences depending on various factors.
+This diversity in our dataset enhances the accuracy and realism of the personalized sticker retrieval task, allowing it to better reflect the range of behaviors found in real-life scenarios.
+Figure 8.
+User D’s consistent usage of the green laughing frog sticker (Pepe frog) from February to January, with occasional months of no sticker usage, highlighting a strong preference for this particular sticker over time.
+5.4. User Behavior with Different Responders
+We highlight two examples of how sticker usage varies with different recipients.
+In the first case, Figure 9, User E replies to both Users F and G using a consistent set of stickers—primarily a nonchalant boy sticker—demonstrating stable preferences regardless of recipient.
+Figure 9. User E consistently uses the red t-shirt with smirk boy sticker and the grey cartoon straight-face sticker when replying to Users F and G
+Figure 10. User H displays contrasting style preferences when replying to Users I and J, using cute, cuddly bear stickers for User I and black-and-white, melancholic stickers for User J.
+Contrastively, User H adapts their sticker style depending on the recipient: using cheerful, colorful stickers with User I, but switching to melancholic, monochromatic stickers with User J, Figure 10.
+In conclusion, these examples demonstrate that our dataset captures meaningful variations in sticker usage based on the context of the interaction.
+This variability, driven by the recipient, intuitively aligns with our commonsense understanding of how individuals communicate differently with different people.
+U-Sticker effectively captures this behavior, which in turn empowers related tasks, such as personalized sticker retrieval, by enabling a deeper and more context-aware analysis of user preferences.
+6. Application Case 2: Sticker Recommendation
+In this section, we evaluate the feasibility of our dataset through quantitative experiments. Specifically, we evaluate U-Sticker dataset on current sticker recommendation.
 6.1. Experiment Settings
-We evaluate two baseline methods for emoticon recommendation, described as follows:
+We evaluate two baseline methods for sticker recommendation, described as follows:
   1. (1)
@@
     SRS (Gao et al., 2021): This approach ranks a list containing negative and one positive sample, returning the ranked results.
-We conduct experiments on three distinct perspectives of the dataset: (1) the English and Chinese subsets, (2) the domain-specific subset, and (3) the complete dataset.
-6.1.1. English and Chinese Subset
-Since MOD can only handle English and Chinese, we select groups with more than 80% of text being either English or Chinese and train the model on it. We do the same for SRS.
-6.1.2. Domain
-To interpretate the performance between different domains, we separate the dataset by domains as shown in Table 3. Since SRS is not limited by multilingualism, we evaluate using SRS.
-6.1.3. Complete
-Finally, the entire dataset is utilized for evaluation on SRS.
+We conduct experiments on three distinct perspectives of the U-Sticker dataset: (1) the English and Chinese subsets, (2) the domain-specific subset, and (3) the complete dataset.
+Since MOD can only handle English and Chinese, we select groups with more than 80% of text being either English or Chinese and train the model on it.
+We do the same for SRS.
+This forms the English and Chinese Subset setting.
+To interpretate the performance between different domains, we separate the dataset by domains as shown in Table 3.
+Since SRS is not limited by multilingualism, we evaluate using SRS.
+This forms the Domain setting.
+Finally, the complete U-Sticker dataset is evaluated on SRS.
 6.2. Results
@@
 For instance, in the 1:9 ratio, the MOD method for English achieves an MRR of 0.289, while the Chinese MOD method only reaches 0.317, this could be due to the smaller data size of Chinese in comparison to English.
-Table 4. emoticon Retrieval Results for MOD and SRS (English and Chinese). We shortform Negative Samples for NS.
+Table 4. Sticker Retrieval Results for MOD and SRS on the English and Chinese. Negative Samples is shortformed NS.
 Lang NS Method MRR R@1 R@3 R@5 R@10
@@
 6.2.2. Domain
-Table 5 reveal that the emoticons domain performs the best overall, with an MRR of 0.335, R@1 of 0.143, R@3 of 0.365, and R@5 of 0.552.
+Table 5 reveal that the Media Sharing domain performs the best overall, with an MRR of 0.335, R@1 of 0.143, R@3 of 0.365, and R@5 of 0.552.
 On the other hand Arts domain has a relatively lower performance, possibly due to the higher sparsity (0.019143).
-Table 5. Results of SRS on individual domains.
+Table 5. Results of SRS on individual domains. Media Sharing domain performs the best overall, whereas Arts domain has a relatively lower performance possibly due to higher sparsity.
 Domain MRR R@1 R@3 R@5
@@
 As the number of negative sample increase, a decline in accuracy is observed, which is consistent with general recommendation models.
-Table 6. Performance metrics for varying negative samples across multiple evaluation criteria on the entire dataset using SRS method.
+Table 6. Performance metrics for varying NS across multiple evaluation criteria on the entire dataset using SRS method.
 NS Evaluation Metrics
@@
 7. Discussions
-7.1. Other Potential Applications
-Beyond user behavior analysis and emoticon recommendation, the dataset presents several opportunities for further research.
-One potential avenue is a more extensive User Behavior Modeling, which could involve analyzing temporal changes in user behavior and investigating whether emoticon usage varies based on the responder.
-Additional potential applications include, but are not limited to, Personalized Emoticon Retrieval and Generative AI, both of which have previously faced dataset limitations (Shen et al., 2024; Xu et al., 2024).
+7.1. Other Potential Applications of U-Sticker
+Beyond user behavior analysis and sticker recommendation, the U-Sticker dataset presents several opportunities for further research.
+One potential avenue is a more extensive User Behavior Modeling, which could involve more quantitative user analysis methods, or involve analyzing temporal changes in user behavior and investigating whether sticker usage varies based on the responder.
+Additional potential applications include, but are not limited to, Personalized Sticker Retrieval and Generative AI, both of which have previously faced dataset limitations (Shen et al., 2024; Xu et al., 2024).
 We acknowledge that there are several limitations in our work.
@@
 8. Conclusion
-In this paper, we introduce our emoticon dataset, the largest emoticon dataset to date, comprising data from 22K users, 370.2K emoticons, and 8.3M conversation messages.
+In this paper, we introduce U-Sticker dataset, the largest sticker dataset to date, comprising data from 22.6k users, 370.2k stickers, and 8.8M conversation messages.
 This multi-domain dataset encompasses a wealth of diverse information, capturing temporal, multilingual, and cross-domain behaviors not found in previous datasets.
-Extensive quantitative and qualitative experiments demonstrate the practical applications of our dataset in user behavior modeling and personalized emoticon recommendation.
+Extensive quantitative and qualitative experiments demonstrate the practical applications of U-Sticker in user behavior modeling and personalized sticker recommendation.
 It also holds potential for further research in areas such as personalized retrieval and conversational studies.